# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

### v3

@@ Abstract
We prove a new asymptotic equipartition property for the perplexity of long texts generated by a language model and present supporting experimental evidence from open-source models. Specifically we show that the logarithmic perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions. This defines a “typical set” that all long synthetic texts generated by a language model must belong to. We show that this typical set is a vanishingly small subset of all possible grammatically correct outputs. These results suggest possible applications to important practical problems such as (a) detecting synthetic AI-generated text, and (b) testing whether a text was used to train a language model. We make no simplifying assumptions (such as stationarity) about the statistics of language model outputs, and therefore our results are directly applicable to practical real-world models without any approximations.
Language Models, Perplexity, Asymptotic Equipartition, Typical Sets
1 Introduction and Motivation.
@@
Consider a generative model, defined as an algorithm that takes a user input ${\mathbf{X}}$ and produces an output ${\mathbf{Y}}$ that statistically resembles data from a natural source. A specific type of generative model is a large language model (LLM) whose output ${\mathbf{Y}}$ is a body of text and the input is a text user prompt ${\mathbf{X}}$. State-of-the-art LLMs (Anthropic, 2024; OpenAI, 2024) are now able to produce detailed and information-rich text outputs such as entire screenplays and book-length manuscripts from a short and simple user prompt. Furthermore, LLMs produce text that can imitate human language well-enough to pass the Turing test (Pinar Saygin et al., 2000) i.e. resembles text created by humans well enough to be convincing to human observers. In this work, we argue that an LLM is strongly constrained by statistical laws. Specifically, we show that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions. This means that any language model is constrained to only output text strings from a typical set, which is an exponentially vanishing subset of all possible grammatically correct strings.
1.1 Contribution: Equipartition Property for Perplexity
Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for language models (Meister & Cotterell, 2021). It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and cross-entropy (Cover & Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg & Keller, 2008; Cohen & Pakhomov, 2020) phenomena at least partially. Many “AI detection” tools for identifying synthetic text from language models are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
@@ 1.1 Contribution: Equipartition Property for Perplexity
Our main result is a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory. The simplest version of the AEP states that a long sequence of independent and identically distributed (iid) random symbols is likely to be a typical sequence (Csiszar, 1998) defined by the property that the empirical distribution of the occurrence of different symbols within the sequence is very close to the distribution from which the symbols were drawn.
The AEP itself can be thought of as a variant of the Law of Large Numbers (LLN) (Idele, 2018) applied to a log likelihood function, and just like the LLN, there is an extensive literature on extending (Moy, 1961) the AEP to a sequence of nearly-independent but not identically distributed random variables (Nishiara & Morita, 2000; Timo et al., 2010). The outputs of non-trivial language models, however, cannot be reasonably modeled as a sequence of nearly-independent symbols (Brown et al., 2020).
@@
Our main contribution is to formulate and provide theoretical and experimental justification for a version of the AEP that applies to real-world practical LLMs without making simplifying assumptions about their statistics. A secondary contribution is to show possible applications of this theory to practical problems such as AI text detection and LLM dataset inference.
2 Background and Definitions
@@
Let M be a generative model described by ${\mathbf{Y}}=m({\mathbf{X}},{\mathbf{W}})$, where the output ${\mathbf{Y}}$ consisting of a (potentially infinite) string of tokens ${\mathbf{Y}}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}~{}\dots]$ is a deterministic function of user prompt ${\mathbf{X}}$ and a pseudo-random sequence ${\mathbf{W}}$. Each token is chosen from a finite set of tokens $Y_{n}\in\mathcal{T}$. The model M can produce a number of different output strings ${\mathbf{Y}}$ for the same user prompt ${\mathbf{X}}$ corresponding to different values of the pseudo-random sequence ${\mathbf{W}}$. This defines a probability distribution $P({\mathbf{Y}}|{\mathbf{X}})$, or more formally, a sequence of probability distributions $P({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ where ${\mathbf{Y}}_{N}\doteq[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ is the substring of ${\mathbf{Y}}$ consisting of the first $N$ tokens.
2.1 Sequential LLM
Practical implementations of LLMs specify the probability distribution iteratively (Radford et al., 2018):
@@ 2.1 Sequential LLM
$P$ $({\mathbf{Y}}_{N}|{\mathbf{X}})=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),$
$~{}\mathrm{where}~{}p_{n}(Y_{n})=P(Y_{n}|Y_{1},~{}Y_{2},~{}Y_{n-1},{\mathbf{X}})$ (2)
Given a string ${\mathbf{Y}}$, open-source LLMs can be programmed to print out the distributions $p_{n}(Y_{n})$ from which its tokens were selected. Specifically, given a user prompt ${\mathbf{X}}$ and a string of tokens ${\mathbf{Y}}_{N}\equiv[Y_{1},~{}Y_{2},~{}\dots,~{}Y_{N}]$, it is possible to get a complete listing of the distributions $p_{n}(Y_{n}),~{}n=1\dots N$. Note that complete knowledge of the $p_{n}(y)$ is not the same as complete knowledge of $P({\mathbf{Y}}|{\mathbf{X}})$ even for a fixed user prompt ${\mathbf{X}}$. As an example, the former requires knowledge of $p_{3}(y)\equiv P(y|Y_{1},Y_{2},{\mathbf{X}})$ for different values of $y\in\mathcal{T}$, but only conditioned on the specific values of $Y_{1},~{}Y_{2}$ contained in one specific string ${\mathbf{Y}}$.
@@
The perplexity $\mathrm{perp}_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a (finite length) text string ${\mathbf{Y}}_{N}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string ${\mathbf{Y}}$. It is usually more convenient to work with the log-perplexity $l_{M}({\mathbf{Y}}_{N})$:
$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(\mathrm{perp}_{M}({\mathbf{Y}}_{N})\right)$
$\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
@@ 2.2 A Toy Problem
Let $\alpha(y),~{}\beta(y)$ be two fixed probability distributions over the (discrete) set $\mathcal{T}$ of tokens. Consider a toy problem involving two language models A and B, that each generate a string of tokens ${\mathbf{Y}}=[Y_{1},~{}Y_{2},~{}\dots]$ where each token is generated iid from the distribution $\alpha(y)$ and $\beta(y)$ respectively. The iid assumption implies that the tokens $Y_{n}$ can be thought of as being generated by a stationary and ergodic random process [^1] Strictly speaking, a stationary random process “starts” at $n=-\infty$ rather than $n=1$; here we ignore the tokens produced by the assumed random process before $n=1$..
Consider a long string ${\mathbf{A}}_{N}\doteq[A_{1},~{}A_{2},~{}\dots,~{}A_{N}],~{}N\gg 1$ randomly generated from model A. Let $p_{\mathbf{A}}(y)$ denote the empirical distribution of $y\in\mathcal{T}$ in the string ${\mathbf{A}}_{N}$:
$p_{\mathbf{A}}(y)$ $\doteq\frac{n_{\mathbf{A}}(y)}{N}\equiv\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}(A_{n}\equiv y),~{}\forall y\in\mathcal{T}$ (4)
@@
$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))$
$\equiv-\sum_{y\in\mathcal{T}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
where $H(\gamma_{1},\gamma_{2})\doteq-\sum_{y\in\mathcal{T}}\gamma_{1}(y)\log_{2}(\gamma_{2}(y))$ is the cross-entropy between two distributions $\gamma_{1}(y),~{}\gamma_{2}(y)$ over $y\in\mathcal{T}$.
It is well-known $H(\gamma_{1},\gamma_{2})\geq H(\gamma_{1})$ with equality when $\gamma_{1}\equiv\gamma_{2}$, where $H(\gamma)\doteq-\sum_{y\in\mathcal{T}}\gamma(y)\log_{2}(\gamma(y))$ is the entropy of distribution $\gamma$. Thus we have from (5):
$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(p_{\mathbf{A}})$
${\mathrm{with~{}equality~{}iff~{}}}$ $p_{\mathbf{A}}(y)\equiv\alpha(y),~{}\forall y\in\mathcal{T}$ (6)
@@
Simple AEP. For a long text string ${\mathbf{A}}_{N}=[A_{1},~{}A_{2},~{}\dots,~{}A_{N}]$ of iid tokens $A_{n}$ drawn from a distribution $\alpha(y)$, the log perplexity $l_{A}({\mathbf{A}}_{N})$ as defined in (3) is close to the entropy of the distribution $H(\alpha)$ of the distribution $\alpha(y)$ with high probability:
$\lim_{N\rightarrow\infty}\Pr[|l_{A}({\mathbf{A}}_{N})-H(\alpha)|>\epsilon]$ $\equiv 0,~{}\forall\epsilon>0$ (7)
Proposition 2.1 can be thought of as a direct consequence of the (weak) Law of Large Numbers [^2] For all of our results in this paper, convergence is in the “weak” sense i.e. convergence in probability. applied to the random variable $l_{M}({\mathbf{A}}_{N})$:
$l_{M}({\mathbf{A}}_{N})$ $\rightarrow\operatorname*{{E}}\left(l_{M}({\mathbf{A}}_{N})\right)\equiv H(\alpha)$ (8)
From (5) and (6), we see that (8) itself is equivalent to $p_{\mathbf{A}}(y)\equiv\frac{n_{\mathbf{A}}(y)}{N}\rightarrow\alpha(y),~{}\forall y\in\mathcal{T}$ i.e. the empirical distribution $p_{\mathbf{A}}(y)$ for long strings ${\mathbf{A}}$ is very close to $\alpha(y)$ with high probability. Putting these observations together we have:
$l_{A}({\mathbf{A}})$ $\equiv H(p_{\mathbf{A}},\alpha)\rightarrow H(p_{\mathbf{A}})$
$\leq H(p_{\mathbf{A}},\beta)\equiv l_{B}({\mathbf{A}})$ (9)
Intuitively, (9) states that a long string ${\mathbf{A}}$ generated from model A is likely to have a lower perplexity for model A than for any other model B. This means that a model trained to have low perplexity over a set of reference texts will generate text strings that are statistically similar to the reference texts. This is the theoretical justification for using perplexity as a loss function for training language models and it is an excellent justification - provided only that we accept the Shannon model of language i.e. the idea that languages can be reasonably modeled as a stochastic sequence of tokens (Jakobson, 1961). This same reasoning also underlies popular tools for LLM-related applications as we discuss in Section 5.

### v4

@@ Abstract
We prove a new asymptotic unequipartition property for the perplexity of long texts generated by a language model and present supporting experimental evidence from open-source models. Specifically we show that the logarithmic perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions. This defines a “typical set” that all long synthetic texts generated by a language model must belong to. We refine the concept of ”typical set” to include only grammatically correct texts. We then show that this refined typical set is a vanishingly small subset of all possible grammatically correct texts for a very general definition of grammar. This means that language models are strongly constrained in the range of their possible behaviors and outputs. We make no simplifying assumptions (such as stationarity) about the statistics of language model outputs, and therefore our results are directly applicable to practical real-world models without any approximations. We discuss possible applications of the typical set concept to problems such as detecting synthetic texts and membership inference in training datasets.
1 Introduction and Background.
@@
Consider a generative model, defined as an algorithm that takes a user input $\mathbf{X}$} and produces an output $\mathbf{Y}$} that statistically resembles data from a natural source. A specific type of generative model is a large language model (LLM) whose output $\mathbf{Y}$} is a body of text and the input is a text user prompt $\mathbf{X}$}. State-of-the-art LLMs (Anthropic, 2024; OpenAI, 2024) are now able to produce detailed and information-rich text outputs such as entire screenplays and book-length manuscripts from a short and simple user prompt. They can also pass the Turing test (Pinar Saygin et al., 2000) i.e. imitate human-written text well enough to be convincing to human observers. In this work, we argue that an LLM is strongly constrained by statistical laws. Specifically, we show that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions. This means that any language model is constrained to only output text strings from a typical set, which we show, is an exponentially vanishing subset of all possible grammatically correct strings. This involves a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory and refining it for language models to account for the concept of grammatical correctness.
This work touches on ideas such as entropy, equipartition theorems and computational linguistics/natural language processing (CL/NLP), each of which boast of a vast body of previous work. We now present a brief literature survey to place our work in context.
1.1 Equipartition Theorems
@@
The simplest version of the AEP states that a long sequence of independent and identically distributed (iid) random symbols is likely to be a typical sequence (Csiszar, 1998) defined by the property that the empirical distribution of the occurrence of different symbols within the sequence is very close to the distribution from which the symbols were drawn. The AEP itself can be thought of as a consequence of the Law of Large Numbers (LLN) (Idele, 2018) that asserts the convergence of a log-likelihood random variable to a limiting “entropy rate” (Algoet & Cover, 1988).
A weak AEP for iid sequences with a finite alphabet was originally introduced by Shannon (1948) followed by strong versions for stationary, ergodic sequences by McMillan (1953) and Breiman (1957), extensions to infinite alphabets by Chung (1961), and continuous-valued sequences in Barron (1985). Thus, the AEP is also commonly referred to as the Shannon-McMillan-Breiman (SMB) Theorem Hamdan et al. (2008). Generalizations of the SMB Theorem (Moy, 1961) in the form of ergodic theorems are now available for many classes of stationary sequences (Pollicott & Yuri, 1998).
Extending the AEP for iid sequences to a non-stationary sequence of independent, but non-identically distributed symbols is straightforward (Tang et al., 2024). AEPs for word-valued functions of stationary sequences are considered in (Nishiara & Morita, 2000; Timo et al., 2010). However, beyond this, the literature on the AEP for non-stationary sequences is much more limited. Closest to our work is the analysis in Yang & Liu (2004) and Wang & Yang (2016) that consider a very general class of non-homogeneous Markov processes, but assume the existence of a unique limiting distribution for the Markov process to establish a constant entropy rate.
Our key point of departure is to not require equal partitions: without stationarity or similarly restrictive assumptions on the statistics of the process, the likelihoods of long sequences do not all become asymptotically equal. Crucially, we show that the the typical sets defined by the resulting Un-Equipartition Property still retain certain essential properties that make the typical set concept so powerful for stationary processes.
1.2 Computational Models of Language
Statistical analysis of languages in its modern form dates back to the work of (Shannon, 1951) on the entropy of written English. These statistical ideas competed (Chomsky, 1959) with structural theories of language (Chomsky, 2014) in the early days of modern linguistics. Early methods such as n-gram models (Manning & Schütze, 1999) evolved into more mathematically sophisticated methods such as Hidden Markov Models (HMMs) (Rabiner, 1989). The statistical approach has now become dominant after the recent introduction of neural network architectures (Bengio et al., 2003), culminating in the development of powerful large language models (LLMs) based on transformer networks (Vaswani et al., 2017). While we know that LLMs can generate complex and coherent text outputs, their representation of language structures and grammar remains a topic of research Reizinger et al. (2024); Mészáros et al. (2024).
Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for training language models (Meister & Cotterell, 2021). It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and entropy (Cover & Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg & Keller, 2008; Cohen & Pakhomov, 2020) phenomena at least partially. Many “AI detection” tools are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
1.3 Contribution: An Un-Equipartition Property for Unnatural Language
@@
Our main contributions are (a) to formulate and provide theoretical and experimental justification for a version of the AEP that applies to real-world practical LLMs without any approximations, and (b) use this AEP to define a concept of typical set specifically adapted for language models. While our focus is on the theoretical development, we also briefly discuss possible applications to practical problems such as AI text detection and LLM dataset inference.
Unlike computational linguistics, we do not study natural language; instead our exclusive focus is on unnatural language i.e. synthetic text created by LMMs. Natural languages are shaped by complex social, psychological and biological processes Pierce (1968) and they can only ever be approximated by computational models Bar-Hillel (1962). In contrast, LLMs are machines whose internals, while complex, are in theory completely knowable.
Thus, it is possible for us to derive exact laws that govern the outputs of LLMs in a way that is not possible to do for natural language. However, this requires us to be disciplined about not making any assumptions or approximations about the statistics of LLM outputs.
1.4 Notation
A sequence of symbols is denoted by boldface, uppercase identifiers e.g. ${\mathbf{A}},~{\mathbf{B}}$. These can be fixed, constant sequences or random sequences. The vector consisting of the first $N$ elements of a sequence $\mathbf{A}$} is denoted by ${\mathbf{A}}_{N}\equiv[A_{1},~A_{2},~\dots,~A_{N}]$. All random variables in this work take discrete-values from a finite set. The probability of an event $E$ is $\Pr(E)$, the distribution or probability mass function of a random variable $A_{m}$ is $p(A_{m})$. Sets are denoted by calligraphic symbols e.g. $\mathcal{A}$. The binary entropy of a distribution $p(A_{m})$ is $H(p(A_{m}))$ [^1] We will always speak of the entropy of distributions and avoid speaking of the entropy of random variables.. The symbol $\equiv$ denotes a simple identity while == denotes a mathematical inference and $\doteq$ a definition.
2 Definitions and Problem Statement
@@
Let M be a generative model whose output $\mathbf{Y}$} consisting of a sequence of tokens ${\mathbf{Y}}=[Y_{1},~Y_{2},~\dots Y_{N}~\dots]$, chosen from a finite token dictionary $Y_{n}\in\mathcal{Y}$, is a stochastic function of user prompt $\mathbf{X}$}. The model M is defined by a probability distribution $p({\mathbf{Y}}|{\mathbf{X}})$, or more formally, a sequence of probability distributions $p({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{Y}^{N}$ where ${\mathbf{Y}}_{N}\doteq[Y_{1},~Y_{2},~\dots Y_{N}]$ is the substring of $\mathbf{Y}$} consisting of the first $N$ tokens. Practical implementations of LLMs specify the probability distribution iteratively (Radford et al., 2018):
@@
$p$ $({\mathbf{Y}}_{N}|{\mathbf{X}})=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),~\mathrm{where}~p_{n}(Y_{n})=p(Y_{n}|Y_{1},~Y_{2},~Y_{n-1},{\mathbf{X}})$ (2)
Given a string $\mathbf{Y}$}, open-source LLMs can be programmed to print out the distributions $p_{n}(Y_{n})$ from which its tokens were selected. Specifically, given a user prompt $\mathbf{X}$} and a string of tokens ${\mathbf{Y}}_{N}\equiv[Y_{1},~Y_{2},~\dots,~Y_{N}]$, it is possible to get a complete listing of the distributions $p_{n}(Y_{n}),~n=1\dots N$.
@@
The perplexity $\mathrm{perp}_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a text string ${\mathbf{Y}}_{N}=[Y_{1},~Y_{2},~\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string $\mathbf{Y}$}. It is usually more convenient to work with the log-perplexity $l_{M}({\mathbf{Y}}_{N})$:
$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(\mathrm{perp}_{M}({\mathbf{Y}}_{N})\right)\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
@@ 2.1 A Toy Problem
Let $\alpha(y),~\beta(y)$ be two fixed probability distributions over the (discrete) set $\mathcal{Y}$ of tokens. Consider a toy problem involving two language models A and B, that each generate a string of tokens ${\mathbf{Y}}=[Y_{1},~Y_{2},~\dots]$ where each token is generated iid from the distribution $\alpha(y)$ and $\beta(y)$ respectively. The iid assumption implies that the tokens $Y_{n}$ can be thought of as being generated by a stationary and ergodic random process [^2] Strictly speaking, a stationary random process “starts” at $n=$-$\infty$ rather than $n=1$; here we ignore the tokens produced by the assumed random process before $n=1$..
Consider a long string ${\mathbf{A}}_{N}\doteq[A_{1},~A_{2},~\dots,~A_{N}],~N\gg 1$ randomly generated from model A. Let $p_{\mathbf{A}}(y)$ denote the empirical distribution of $y\in\mathcal{Y}$ in the string ${\mathbf{A}}_{N}$:
$p_{\mathbf{A}}(y)$ $\doteq\frac{n_{\mathbf{A}}(y)}{N}\equiv\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}(A_{n}\equiv y),~\forall y\in\mathcal{Y}$ (4)
@@
$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))\equiv-\sum_{y\in\mathcal{Y}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
where $H(\gamma_{1},\gamma_{2})\doteq-\sum_{y\in\mathcal{Y}}\gamma_{1}(y)\log_{2}(\gamma_{2}(y))$ is the cross-entropy between distributions $\gamma_{1}(y),~\gamma_{2}(y)$ over $y\in\mathcal{Y}$. It is well-known $H(\gamma_{1},\gamma_{2})\geq H(\gamma_{1})$ with equality when $\gamma_{1}\equiv\gamma_{2}$, where $H(\gamma)\doteq-\sum_{y\in\mathcal{Y}}\gamma(y)\log_{2}(\gamma(y))$ is the entropy of distribution $\gamma$. Thus we have from (5):
$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(p_{\mathbf{A}}){\mathrm{with~equality~iff~}}p_{\mathbf{A}}(y)\equiv\alpha(y),~\forall y\in\mathcal{Y}$ (6)
@@
A Simple Cross-AEP. For a long text string ${\mathbf{A}}_{N}=[A_{1},~A_{2},~\dots,~A_{N}]$ of iid tokens $A_{n}$ drawn from a distribution $\alpha(y)$, the log perplexity $l_{B}({\mathbf{A}}_{N})$ for a model B as defined in (3) almost always approaches the cross-entropy of the distributions $H(\alpha,\beta)$:
$\lim_{N\to\infty}l_{B}({\mathbf{A}}_{N})$ $\equiv H(\alpha,\beta),~\mathrm{and}~\lim_{N\to\infty}l_{A}({\mathbf{A}}_{N})\equiv H(\alpha)$ (7)
From (5) and (6), we see that (7) itself is equivalent to $p_{\mathbf{A}}(y)\equiv\frac{n_{\mathbf{A}}(y)}{N}\rightarrow\alpha(y),~\forall y\in\mathcal{Y}$ i.e. the empirical distribution $p_{\mathbf{A}}(y)$ for long strings $\mathbf{A}$} almost always converges to $\alpha(y)$. Putting these observations together we have:
$\lim_{N\rightarrow\infty}l_{A}({\mathbf{A}}_{N})$ $\equiv H(\alpha)\leq H(\alpha,\beta)\equiv\lim_{N\rightarrow\infty}l_{B}({\mathbf{A}}_{N})$ (8)
Intuitively, (8) states that a long string ${\mathbf{A}}_{N}$ generated from model A is likely to have a lower perplexity for model A than for any other model B. This means that a model trained to have low perplexity over a set of reference texts will generate text strings that are statistically similar to the reference texts. This is the theoretical justification for using perplexity as a loss function for training language models and it is an excellent justification - provided only that we accept the Shannon model of language i.e. the idea that languages can be reasonably modeled as a stochastic sequence of tokens (Jakobson, 1961). Testing for abnormally high or low perplexity is an important tool for LLM applications as we discuss in Section 6.